Solve
a) (x^2 + 3x - 3) ^ 1/2 = x+1
Can you please help me out? Thanks so much in advance:)

Answer by stanbon(75887)

Solve
a) (x^2 + 3x - 3) ^ 1/2 = x+1
-------
Square both sides to get:
x^2+3x-3 = x^2+2x+1
----------
x = 4
================
Cheers,
Stan H.